Exponential equations without using logs. Solve for x:
3^(4x+1)= 27

Answer by stanbon(75887)

Exponential equations without using logs. Solve for x:
3^(4x+1)= 27
3^(4x+1) = 3^3
Since the bases are the same, the exponents are equal.
EQUATION:
4x+1 = 3
4x = 2
x = 1/2
============
Cheers,
Stan H.